OpenStudy (anonymous):

(-3x^-6)(7x^-7)= -21/x^13 true/false?

OpenStudy (aripotta):

false

OpenStudy (aripotta):

i believe it would be -21/x^-13

OpenStudy (aripotta):

since when you multiply, you add the exponents. -6 + (-7) = -6 - 7 = -13

OpenStudy (aripotta):

erm wait.

OpenStudy (aripotta):

does that slash mean divide?

OpenStudy (anonymous):

well the problem in the book also adds to simplify using positive exponents?

OpenStudy (anonymous):

yeah.

OpenStudy (aripotta):

i still think it's false

OpenStudy (aripotta):

idk for sure. you should ask someone else